The density of neon will be highest at
a. STP
b. 0° C, 2 atm
c. 273° C, 1 atm
d. 273° C, 2 atm

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. The ideal gas equation : pV = nRT = w/M RT ⇒ pM = (w/V) RT = dRT (d = density) ⇒ d = pM/RT ie, density will be greater at low temperature and high pressure. 14. The rate of diffusion of methane at a given temperature is twice that of a gas X. The molecular weight of X is a. 64.0 b. 32.0 c. 4.0 d. 8.0 Sol. = 2 = √(M_X/16) ⇒ MX = 64